need help in business math application of portion, base, and rate formulas
assume that in the year 2010, 800,000 people attended the Christmas eve celebration at Walt Disney World. if in 2011, attendance for the Christmas eve celebration is expected to increase by 35%. what is the total number of people expected at Walt Disney World for this event?

Answer by checkley79(3341)

800,00+.35*800,000=800,000+280,000=1,080,000 PEOPLE ANS.